OpenStudy (anonymous):

i am working on high school geometry, i have to find the lateral and total surface areas of a cylinder with a base area of 169(pi) and a height that is 2 times the radius, i have been trying to solve the base area equation by setting it equal to 2(pi)raduis squared but i am stuck... help pleassseee?

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ok so the area of the base is 169pi

OpenStudy (anonymous):

so what do they mean by base area?

OpenStudy (anonymous):

do they mean the area of the circle at the bottom?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well the bottom of a cylinder is a circle

OpenStudy (anonymous):

the area of a circle is pi*r^2

OpenStudy (anonymous):

so pi*r^2 = 169pi

OpenStudy (anonymous):

r^2 = 169pi/pi = 169

OpenStudy (anonymous):

r = sqrt(169) = 13

OpenStudy (anonymous):

if the height is 2*r, then the height is 2*13 = 26

OpenStudy (anonymous):

you should be able to get the total surface area, volume, etc... from there

OpenStudy (anonymous):

does that make sense?

OpenStudy (preetha):

Very smart, superduper

OpenStudy (anonymous):

yes! thank you so much!

OpenStudy (anonymous):

yay! thanks preetha